OpenStudy (anonymous):

what is a quadratic equation for the variable x using -sqrt6, 8sqrt6

OpenStudy (anonymous):

\[-\sqrt{6},8\sqrt{6}\]

OpenStudy (anonymous):

type in ax^2+bx+c=0

OpenStudy (nenadmatematika):